# Semantic Equitable Clustering: A Simple, Fast and Effective Strategy for Vision Transformer

### v1

@@ 4.3 Semantic Segmentation
MPViT-B [28] 105 1186 50.3
CSWin-S [12] 65 1027 50.4
SecViT-B 84 1046 52.0
Swin-B [36] 121 1188 48.1
GC ViT-B [20] 125 1348 49.2
DAT-B [54] 121 1212 49.4
InternImage-B [50] 128 1185 50.8
CSWin-B [12] 109 1222 51.1
SecViT-L 130 1256 53.6
@@
Results.
@@
The results of semantic segmentation can be found in the Tab. 5. All the FLOPs are measured with the input resolution of $512\times 2048$, except the group of the SecViT-T, which are measured with the input resolution of $512\times 512$. SecViT achieves the best performance in all settings. Specifically, with the framework of Sementic FPN, our SecViT-S outperforms FAT-B3 for +0.6 mIoU. And SecViT-B surpasses CSWin-S by a large margin (+1.5 mIoU). For the framework of Upernet, SecViT-S surpasses the recent SMT-S by +1.4 mIoU. SecViT-B and SecViT-L also surpass all their counterparts.
@@ 4.4 Ablation Study
Comparison with Swin-Transformer.
As mentioned earlier, based on the Swin-T architecture, we directly replaced WSA/SWSA with our SECA. As shown in Tab. 6, this resulted in a significant performance improvement (classification accuracy increased by 2.3%).
Comparison with CSWin-Transformer.
As shown in Tab. 6, based on the SecViT-T architecture, we replace SECA with CSWSA from CSwin-Transformer. This leads to a significant drop in classification performance (a decrease of 1.2%). This fully demonstrates the effectiveness of our proposed token clustering method.
LCE.
LCE is a simple depth-wise convolution module used to further enhance the model’s ability to model local information. We conduct ablation experiments on LCE and find that it contributes to a certain improvement in the model’s performance. The results are shown in Tab. 6. Specifically, it boosts the model’s classification accuracy by +0.2.
Model Params(M) FLOPs(G) Top1-acc(%)
Swin-T 29 4.5 81.3
SEC-Swin-T 29 4.9 83.6(+2.3)
SecViT-T 15 2.5 82.6
$\xrightarrow{}$ [12] 15 2.5 81.4(-1.2)
w/o LCE 15 2.5 82.4
w/o CPE 15 2.5 82.4
w/o Conv Stem 14 2.1 82.3
Table 6: Ablation Study.
CPE.
The CPE proposed in CPVT [7] is a plug-and-play, highly portable positional encoding method, widely adopted by many methods to provide positional information to the model. It consists of a residual block containing only a 3x3 depth-wise convolution. Ablation results on CPE in Tab. 6 indicate that it offers some performance gains to the model (approximately +0.2 increase in classification accuracy).
num FLOPs(G) Top1-acc(%)
64, 16, 4, – 4.3 84.0
64, 16, 2, – 4.3 84.1
64, 8, 2, – 4.3 84.1
32, 8, 2, – 4.4 84.2
Table 7: Effect of number of clusters.
Conv Stem.
The Conv Stem is utilized in the early stages of the model to extract more refined local features. As indicated by the data in Tab. 6, the Conv Stem enhances the performance of the model to some extent in classification task. Specifically, the Conv Stem improves the classification accuracy by +0.3.
Effect of number of clusters.
In Tab. 7, we validate the impact of different cluster numbers chosen at each of the first three stages of the model on performance. For SecViT-S, we select 32, 8, and 2 clusters for the first three stages, respectively. This configuration achieves a classification accuracy of 84.2%. As we gradually increase the number of clusters at each stage, the model’s overall computation decreases, with a slight drop in accuracy. This demonstrates the robustness of Semantic Clustering with respect to the number of clusters.
4.5 Visualization of Semantic Clustering
Figure 4: Visualization results for SecViT-B. The first row shows the original images, the second row displays the clustering results of the first stage (32 clusters), the second row displays the clustering results of the second stage (8 clusters), and the third row shows the clustering results of the third stage (2 clusters).
To further understand how our semantic equitable clustering works, we visualized the clustering results of some samples. As shown in Fig. 4, we selected the first, second, and third stages of the pre-trained SecViT-B to observe the clustering results of different tokens. In the figure, patches with the same color belong to the same cluster. It is evident that in the first and second stages, our model tends to capture detailed information, while in the third stage, it tends to capture semantic information.
@@ 5 Conclusion
We propose a simple and straightforward clustering method for Vision Transformers—Semantic Equitable Clustering (SEC). This method assigns each token to a cluster by calculating the similarity between each token and a global token, then sorting them accordingly. Our clustering method takes into account the semantic information contained in the tokens, completes the clustering in a single step, and ensures an equal number of tokens in each cluster, facilitating efficient parallel processing on modern GPUs. Based on Semantic Equitable Clustering, we designed SecViT, a versatile vision backbone that achieves impressive results across various vision tasks, including image classification, object detection, instance segmentation, and semantic segmentation.
Supplementary Material
@@
SecViT-S [3, 4, 18, 4] [64, 128, 256, 512] [2, 4, 8, 16] 3 27 4.4
SecViT-B [4, 8, 26, 8] [80, 160, 320, 512] [5, 5, 10, 16] 3 55 9.7
SecViT-L [4, 8, 26, 8] [112, 224, 448, 640] [7, 7, 14, 20] 3 96 18.1
SEC-Swin-T [2, 2, 6, 2] [96, 192, 384, 768] [3, 6, 12, 24] 4 29 4.9
SEC-Swin-S [2, 2, 18, 2] [96, 192, 384, 768] [3, 6, 12, 24] 4 50 9.3
SEC-FasterViT-0 [2, 3, 6, 5] [64, 128, 256, 512] [–, –, 8, 16] 4 25 3.6
SEC-FasterViT-1 [1, 3, 8, 5] [80, 160, 320, 640] [–, –, 8, 16] 4 41 5.6
@@
Appendix A Architecture Details
SecViT’s architecture details are illustrated in Table 8. In SecViT, we adopt four $3\times 3$ convolutions to embed the input image into tokens, batch normalization and GELU are used after each convolution. $3\times 3$ convolutions with stride 2 are used between stages to reduce the feature resolution. $3\times 3$ DWConvs are adopted in CPE. While $5\times 5$ depth-wise convolutions are adopted for LCE.
For SEC-Swin, we adhere to the design principles of the Swin-Transformer [36] without using additional structures such as CPE or Conv Stem.
For SEC-FasterViT, we adhere to the design principles of FasterViT [19]. We use convolutional for fast inference speed in the first two stages of the model and use our Semantic Equitable Clustering based Self-Attention (SECA) in the latter two stages.
For all models, we set the number of clusters in the first three stages to 32, 8, and 2, respectively.
Appendix B Experimental Settings
@@
We adopt the training strategy proposed in DeiT [47], with the only supervision is cross entropy loss. All of our models are trained from scratch for 300 epochs with the input resolution of $224\times 224$. The AdamW is used with a cosine decay learning rate scheduler and 5 epochs of linear warm-up. The batch-size is set to 1024, respectively. We apply the same data augmentation and regularization used in DeiT [47], including RandAugment [9] (randm9-mstd0.5-inc1) , Mixup [60] (prob = 0.8), CutMix [59] (prob = 1.0), Random Erasing (prob = 0.25), and Exponential Moving Average (EMA) [41]. The maximum rates of increasing stochastic depth [23] are set to 0.1/0.15/0.4/0.5 for SecViT-T/S/B/L.
@@
We apply RetinaNet [32], Mask-RCNN [21], and Cascaded Mask R-CNN [3] as the detection frameworks based on the MMDetection [6]. All of our models are trained with “1 $\times$" (12 training epochs) and “3 $\times$ +MS" (36 training epochs with multi-scale training) settings. For the “1 $\times$" setting, images are resized to the shorter side of 800 pixels while the longer side is within 1333 pixels. For the “3 $\times$ +MS", multi-scale training strategy is used to randomly resize the shorter side of images between 480 to 800 pixels. For both frameworks, we use the AdamW with the initial learning rate of 1e-4. For RetinaNet, we set the weight decay to 1e-4. While for Mask-RCNN and Cascaded Mask R-CNN, we set it to 5e-2.
ADE20K Semantic Segmentation.
@@
Appendix C Limitations and Future Work
While the proposed Semantic Equitable Clustering have demonstrated significant efficiency for clustering tokens and impressive performance on many vision tasks, there are still several limitations that need to be addressed in our future work. First, the computational constraints prevented us from experimenting with larger models and datasets such as ImageNet-21k. Exploring the potential of Semantic Equitable Clustering on large-scale datasets and larger model could provide further insights into its scalability and efficiency. In the future, we will strive to validate the performance of our SecViT on large datasets and with larger models.
Appendix D Broader Impact Statement
The development of the Semantic Equitable Clustering (SEC) has the potential to impact the field of computer vision by offering a efficient and semantic-related approach to group tokens in Vision Transformer. By clustering the tokens, SEC reduces computational load and improves the efficiency of vision models, which could lead to broader applications.
The proposed SecViT is a general vision backbone that can be applied on different vision tasks, e.g., image classification, object detection, instance segmentation, and semantic segmentation. It has no direct negative social impact. Possible malicious uses of SecViT as a general-purpose backbone are beyond the scope of our study to discuss.
Appendix E Code
We provide the code of our Semantic Equitable Clustering base Self-Attention.
1class SECA(nn.Module):
2
3 def __init__(self, embed_dim, num_heads, num_clusters):
4 super().__init__()
5 self.num_clusters = num_clusters
6 self.embed_dim = embed_dim
7 self.num_heads = num_heads
8 self.head_dim = embed_dim // num_heads
9 self.scaling = self.head_dim ** -0.5
10 self.qkv = nn.Conv2d(embed_dim, embed_dim*3, 1, bias=True)
11 self.out_proj = nn.Conv2d(embed_dim, embed_dim, 1, bias=True)
12
13 def sort_sim(self, q: torch.Tensor, k: torch.Tensor):
14 ’’
15 q: b n (h w) d
16 k: b n (h w) d
17 ’’
18 cls_token = q.mean(2, keepdim=True) # (b n 1 d)
19 cls_token = F.normalize(cls_token, dim=-1)
20 k = F.normalize(k, dim=-1)
21 cos_sim = cls_token @ k.transpose(-1, -2) # (b n 1 (h w))
22 sort_indices = torch.argsort(cos_sim, dim=-1)
23 return sort_indices.squeeze(2) # (b n (h w))
24
25 def sort_tensor(self, indices: torch.Tensor, target: torch.Tensor):
26 ’’
27 indices: (b n l)
28 target: (b n l d)
29 ’’
30 sorted_tensor = torch.gather(target, dim=-2, index=indices.unsqueeze(-1).expand_as(target))
31 return sorted_tensor
32
33 def restore_tensor(self, indices: torch.Tensor, target: torch.Tensor):
34 ’’
35 indices: (b n l)
36 target: (b n l d)
37 ’’
38 restored_tensor = torch.gather(target, dim=-2, index=indices.sort(dim=-1).indices.unsqueeze(-1).expand_as(target))
39 return restored_tensor
40
41 def forward(self, x: torch.Tensor):
42 ’’
43 x: (b c h w)
44 ’’
45 bsz, _, h, w = x.size()
46 l = h * w
47 assert l % self.num_clusters == 0
48 qkv = self.qkv(x) # (b 3*c h w)
49
50 q, k, v = rearrange(qkv, ’b (m n d) h w -> m b n (h w) d’, m=3, n=self.num_heads)
51 # q, k, v -> b n (h w) d
52 k = k * self.scaling
53
54 if self.num_clusters > 1:
55 indices = self.sort_sim(k, k) # (b n (h w))
56 q = self.sort_tensor(indices, q) # (b n (h w) d)
57 k = self.sort_tensor(indices, k) # (b n (h w) d)
58 v = self.sort_tensor(indices, v) # (b n (h w) d)
59
60 q_l = rearrange(q, ’b n (c l) d -> b n c l d’, c=self.num_clusters)
61 k_l = rearrange(k, ’b n (c l) d -> b n c l d’, c=self.num_clusters)
62 v = rearrange(v, ’b n (c l) d -> b n c l d’, c=self.num_clusters)
63 attn = torch.softmax(q_l @ k_l.transpose(-1, -2), dim=-1) # (b n c l l)
64 v = attn @ v # (b n c l d)
65
66 v = rearrange(v, ’b n c l d -> b n (c l) d’)
67 v = self.restore_tensor(indices, v) # (b n (h w) d)
68 v = rearrange(v, ’b n (h w) d -> b (n d) h w’, h=h, w=w)
69 else:
70 attn = torch.softmax(q @ k.transpose(-1, -2), dim=-1) # (b n (h w) (h w))
71 v = attn @ v # (b n (h w) d)
72 v = rearrange(v, ’b n (h w) d -> b (n d) h w’, h=h, w=w)
73 return self.out_proj(res)

### v2

@@ 4.1 SEC for vision models
StructViT-B [28] 54 529 48.5
RMT-B [15] 57 294 50.4 83 1051 52.0
SECViT-B 60 291 50.7 86 1048 52.2
MogaNet-L [34] 113 1176 50.9
CSWin-B [12] 81 464 49.9 109 1222 51.1
MLLA-B [20] 128 1183 51.9
SGFormer-B [17] 81 475 50.6 109 1304 52.0
RMT-L [15] 98 482 51.4 125 1241 52.8
SECViT-L 103 475 52.2 131 1256 53.8
@@
Model Connector V-T Num Time Speed TextVQA GQA VQAv2 POPE MME
LLaVA-1.5 MLP 576+1 194s 1.0×1.0$\times$1.0 × 58.2 62.0 78.5 86.1 1510.7
LLaVA-1.5+Resampler Resampler 288+1 126s 1.5×1.5$\times$1.5 × 52.1 56.8 76.0 83.1 1393.2
LLaVA-1.5+EViT MLP+EViT 288+1 126s 1.5×1.5$\times$1.5 × 54.6 60.0 77.9 84.3 1483.2
LLaVA-1.5+SEC MLP+SEC 288+1 126s 1.5×1.5$\times$1.5 × 60.1 63.5 78.9 87.7 1510.7
LLaVA-1.5+Resampler Resampler 256+1 116s 1.7×1.7$\times$1.7 × 51.6 56.0 75.2 82.7 1387.2
LLaVA-1.5+Pool MLP+Pool 256+1 116s 1.7×1.7$\times$1.7 × 52.4 57.6 76.4 83.3 1415.5
LLaVA-1.5+EViT MLP+EViT 256+1 116s 1.7×1.7$\times$1.7 × 52.8 59.6 77.1 83.7 1443.7
LLaVA-1.5+SEC MLP+SEC 256+1 116s 1.7×1.7$\times$1.7 × 59.6 63.2 78.6 87.1 1505.2
LLaVA-1.5+Resampler Resampler 192+1 102s 1.9×1.9$\times$1.9 × 50.1 55.2 74.3 82.7 1337.6
LLaVA-1.5+EViT MLP+EViT 192+1 102s 1.9×1.9$\times$1.9 × 51.6 58.6 76.3 83.1 1427.6
LLaVA-1.5+SEC MLP+SEC 192+1 102s 1.9×1.9$\times$1.9 × 57.7 62.7 78.4 86.7 1500.1
LLaVA-1.5+Resampler Resampler 144+1 94s 2.1×2.1$\times$2.1 × 47.6 54.6 72.0 81.9 1293.7
LLaVA-1.5+Pool MLP+Pool 144+1 94s 2.1×2.1$\times$2.1 × 50.0 56.2 73.6 81.9 1310.7
LLaVA-1.5+EViT MLP+EViT 144+1 94s 2.1×2.1$\times$2.1 × 51.2 58.0 76.0 83.1 1393.6
LLaVA-1.5+SEC MLP+SEC 144+1 94s 2.1×2.1$\times$2.1 × 56.8 62.0 78.0 86.1 1487.1
Table 8: Comparison of different vision language connectors on LLaVA-1.5. “V-T Num” denotes the quantity of visual tokens. The computation expense is impacted by V-T Num, with larger values resulting in higher costs. “Speed” refers to the comparative inference velocity relative to LLaVA-1.5. “Time” is the average inference time. Inference speed are measured on the A100.
Model LLM Connector V-T Num Res TextVQA GQA VQAv2 VisWiz SQAimg Speed ($\uparrow$)
7B LLM
Shikra [5] Vicuna-7B MLP 257 224 - - 77.4 - - -
IDEFICS-9B [30] LLaMA-7B Cross Attn 257 224 - 38.4 50.9 35.5 - -
Qwen-VL [1] Qwen-7B Resampler 256 448 - 59.3 78.8 35.2 67.1 -
Qwen-VL-Chat [1] Qwen-7B Resampler 256 448 - 57.5 78.2 38.9 68.2 -
LLaVA-1.5 [38] Vicuna-7B MLP 577 336 58.2 62.0 78.5 50.0 66.8 1.0×1.0$\times$1.0 ×
LLaVA-1.5+SEC (ours) Vicuna-7B MLP+SEC 257 336 59.6 63.2 78.9 52.8 69.6 $\mathbf{1.7\times}$
13B LLM
InstructBLIP [9] Vicuna-13B Q-Former 32 224 - 49.5 - 33.4 63.1 -
BLIP-2 [32] Vicuna-13B Q-Former 32 224 - 41.0 41.0 19.5 61.0 -
LLaVA-1.5 [38] Vicuna-13B MLP 577 336 61.2 63.3 80.0 53.6 71.6 1.0×1.0$\times$1.0 ×
LLaVA1.5+SEC (ours) Vicuna-13B MLP+SEC 257 336 62.3 64.3 80.0 54.7 72.0 $\mathbf{1.8\times}$
Table 9: Results on General VQA tasks.
Model LLM Connector V-T Num Res POPE MMB MM-Vet Speed ($\uparrow$)
7B LLM
MiniGPT-4 [65] Vicuna-7B Resampler 32 224 72.2 24.3 22.1 -
mPLUG-Owl2 [60] LLaMA2-7B Resampler 32 224 - 49.4 - -
LLaMA-AdapterV2 [16] LLaMA2-7B LLaMA-Adapter 257 224 - 41.0 31.4 -
Shikra [5] Vicuna-7B MLP 257 224 - 58.8 - -
Qwen-VL [1] Qwen-7B Resampler 256 448 - 38.2 - -
Qwen-VL-Chat [1] Qwen-7B Resampler 256 448 - 60.6 - -
LLaVA-1.5 [38] Vicuna-7B MLP 577 336 86.1 64.3 31.1 1.0×1.0$\times$1.0 ×
LLaVA1.5+SEC (ours) Vicuna-7B MLP+SEC 145 336 86.1 68.4 31.7 $\mathbf{2.1\times}$
13B LLM
MiniGPT-4 [65] Vicuna-13B Resampler 32 224 - - 24.4 -
BLIP-2 [32] Vicuna-13B Q-Former 32 224 85.3 - 22.4 -
LLaVA-1.5 [38] Vicuna-13B MLP 577 336 86.2 67.7 36.1 1.0×1.0$\times$1.0 ×
LLaVA-1.5+SEC (ours) Vicuna-13B MLP+SEC 145 336 86.4 69.2 37.3 $\mathbf{2.2\times}$
Table 10: Results on benchmark designed for MLLMs.
Strict Comparison with Baselines.
We select two baselines: hierarchical backbone Swin-Transformer [41] and plain backbone DeiT [46] to make a comparison with our SEC based model. In the comparison models (SEC-Swin and SEC-DeiT), we merely substitute the attention mechanism in the original model with our SEC based Self-Attention and without introducing any other modules. As shown in Tab. 2, we conduct experiments on image classification, object detection, insatance segmentation and semantic segmentation, the simple replacement of the attention mechanism yields significant advantages in both performance and efficiency.
In addition to the supervised scenario, we also train the model with SimMIM [56] in the self-supervised scenario. As shown in Tab. 3, SEC also performs exceptionally well in the self-supervised scenario.
Image Classification.
We compare our SECViT with numerous state-of-the-art models, the results are shown in Tab.4. We adopt the training strategy proposed in DeiT [46], with the only supervision is cross entropy loss. All of our models are trained from scratch for 300 epochs with the input resolution of $224\times 224$. SECViT consistently outperforms preceding models across all scales. Notably, SECViT-S attains a Top1-accuracy of 84.3% with a mere 27M parameters and 4.6G FLOPs. For larger models, SECViT-XL achieves a Top1-accuracy of 86.3% with 205M parameters and 36.4G FLOPs. The comparison of the models’ efficiency can be found in Appendix.
Object Detection and Instance Segmentation.
We utilize MMDetection [4] to implement Mask-RCNN [23], Cascade Mask R-CNN [2], and RetinaNet [36] to evaluate the performance of the SECViT. Tab. 6 and Tab. 5 show the results of SECViT with different detection frameworks. The results show that SECViT performs better than its counterparts in all comparisons.
Semantic Segmentation.
@@
We utilize Semantic FPN [29] and UperNet [55] to validate our SECViT’s performance, implementing these frameworks via MMSegmentation [7]. The results of semantic segmentation can be found in the Tab. 7. All the FLOPs are measured with the input resolution of $512\times 2048$. SECViT achieves the best performance in all settings.
4.2 SEC for MLLM
SEC can greatly facilitate the design of vision language connectors in MLLMs. First, we conduct a rigorous comparison between SEC and various baseline vision language connectors based on LLaVA-1.5. Then, we compare LLaVA-1.5+SEC with several popular contemporary MLLMs.
Strict Comparison with Baselines.
In Tab. 8, we strictly compare various commonly used vision language connectors, including MLP, Resampler [1], Pooling, and EViT [35], which has achieved success in the design of ViT. Among these, MLP is the original design in LLaVA-1.5 [39], capable of achieving good results. However, it incurs significant computational cost due to the excessive vision tokens. To address this issue, some connectors attempt to use fewer vision tokens to accelerate LLaVA-1.5. Nonetheless, these adjustments inevitably lead to performance degradation. The results in Tab. 8 show that using SEC can effectively accelerate the inference of LLaVA-1.5 without causing performance degradation, and can even improve the performance of LLaVA-1.5 to a certain extent.
Comparison with Popular MLLMs.
In Tab. 9 and Tab. 10, we compare LLaVA-1.5 equipped with SEC as a vision-language connector with other MLLMs. It is evident that SEC not only enhances the performance of MLLMs across various benchmarks but also significantly improves the efficiency of the models. This fully demonstrates the effectiveness of SEC in extracting visual information.
4.3 Visualization of SEC
To further understand the working mechanism of SEC, we visualize some clustering results for SECViT. As shown in Fig. 5, the left side presents the clustering results of vision tokens at different stages of the model. From the clustering results, we analyze that in the shallow layers, the model distinguishes fine-grained features well, while in the deeper layers, it captures global semantic features effectively. The right side shows the Grad-CAM diagrams at different stages of the model, from which we can draw similar conclusions to the clustering results. More visualization results can be found in Appendix.
Figure 5: Visualization for SEC.
@@ 4.4 Ablation Study
In this section, we present some of the ablation study results for SEC, and more results can be found in the Appendix.
Number of Vision Tokens in Each Clusters.
The number of vision tokens has a significant impact on the performance and speed of the model. We thoroughly investigate the effect of the number of vision tokens on SECViT. As shown in Tab. 11, the number of vision tokens in each cluster greatly influences the model’s performance. Specifically, in downstream dense prediction tasks, having too few tokens in each cluster leads to substantial performance degradation. When the number of tokens in each cluster is too large, the model’s performance does not see a significant improvement, but its speed decreases.
V-T num Params (M) FLOPs (G) Throughput (imgs/s) Acc $AP^{b}$ $AP^{m}$ mIoU
98 15 2.5 2004 82.7 47.8 43.0 47.2
196 15 3.1 1722 83.0 (+0.3) 48.2 (+0.4) 43.4 (+0.4) 47.5 (+0.3)
64 15 2.5 1946 82.7 (+0.0) 47.8 (+0.0) 42.8 (-0.2) 46.9 (-0.3)
49 15 2.4 2102 82.6 (-0.1) 47.5 (-0.3) 42.7 (-0.3) 47.7 (-0.5)
24 15 2.3 2186 82.0 (-0.7) 45.9 (-1.9) 40.6 (-2.4) 44.6 (-2.6)
Table 11: Effect of the number of vision tokens in each cluster. “V-T num” means the number of vision tokens in each cluster. The experiments are conducted based on SECViT-T.
@@ 5 Conclusion
We propose a simple and straightforward clustering method for vision tokens—Semantic Equitable Clustering (SEC). This method assigns each token to a cluster by calculating the similarity between each token and a global token, and completes the whole clustering process in only one step. Our clustering method takes into account the semantic information contained in the tokens, and ensures an equal number of tokens in each cluster, facilitating efficient parallel processing on modern GPUs. Based on Semantic Equitable Clustering, we designed SECViT, a versatile vision backbone that achieves impressive results across various vision tasks, including image classification, object detection, instance segmentation, and semantic segmentation. Besides, SEC can also be conveniently applied to multimodal large language models (MLLM) to serve as a vision language connector and benefits the model’s efficiency.
Appendix A Experimental Details
SECViT’s Architectures.
SECViT’s architecture details are illustrated in Table 12. In SECViT, we adopt four $3\times 3$ convolutions to embed the input image into tokens, batch normalization and GELU are used after each convolution. $3\times 3$ convolutions with stride 2 are used between stages to reduce the feature resolution. $3\times 3$ DWConvs are adopted in CPE. For all models, we set the number of clusters in the first three stages to 32, 8, and 2, respectively.
@@ Appendix A Experimental Details
SECViT-S [4, 4, 18, 4] [64, 128, 256, 512] [2, 4, 8, 16] 3 27 4.6
SECViT-B [4, 8, 26, 9] [80, 160, 320, 512] [2, 4, 8, 16] 3 57 9.8
SECViT-L [4, 8, 26, 9] [112, 224, 448, 640] [4, 8, 14, 20] 3 101 18.2
SECViT-XL [6, 12, 28, 12] [128, 256, 512, 1024] [4, 8, 16, 32] 3 205 36.4
@@
We adopt the training strategy proposed in DeiT [46], with the only supervision is cross entropy loss. All of our models are trained from scratch for 300 epochs with the input resolution of $224\times 224$. The AdamW is used with a cosine decay learning rate scheduler and 5 epochs of linear warm-up. The batch-size is set to 1024, respectively. We apply the same data augmentation and regularization used in DeiT [46], including RandAugment [8] (randm9-mstd0.5-inc1) , Mixup [64] (prob = 0.8), CutMix [62] (prob = 1.0), Random Erasing (prob = 0.25). The maximum rates of increasing stochastic depth [24] are set to 0.1/0.15/0.4/0.5/0.65 for SECViT-T/S/B/L/XL.
@@
We apply RetinaNet [36], Mask-RCNN [23], and Cascaded Mask R-CNN [2] as the detection frameworks based on the MMDetection [4]. All of our models are trained with “1 $\times$” (12 training epochs) and “3 $\times$ +MS” (36 training epochs with multi-scale training) settings. For the “1 $\times$” setting, images are resized to the shorter side of 800 pixels while the longer side is within 1333 pixels. For the “3 $\times$ +MS”, multi-scale training strategy is used to randomly resize the shorter side of images between 480 to 800 pixels. For both frameworks, we use the AdamW with the initial learning rate of 1e-4. For RetinaNet, we set the weight decay to 1e-4. While for Mask-RCNN and Cascaded Mask R-CNN, we set it to 5e-2.
ADE20K Semantic Segmentaion.
@@
Appendix B More experimental Results
Efficiency Comparison.
In Tab. 13, we compare the inference efficiency of various models in detail. From this, we can see that the ViT based on SEC demonstrates the best performance-speed tradeoff.
Model Params(M) FLOPs(G) Throughput(imgs/s) Top1-Acc(%)
DeiT-S [46] 22 4.6 3204 79.8
EViT-DeiT-S (keeprate=0.9) [35] 22 4.0 3428 79.8
SEC-DeiT-S (num_cluster=4) 22 4.1 3412 80.5
DeiT-B [46] 86 17.6 1502 81.8
SEC-DeiT-B 86 14.8 1682 82.4
PVTv2-b1 [49] 13 2.1 2204 78.7
TCFormer-light [63] 14 3.8 417 79.4
MPViT-XS [31] 11 2.9 1496 80.9
BiFormer-T [66] 13 2.2 1634 81.4
CMT-XS [18] 15 1.5 1476 81.8
GC-ViT-XT [22] 20 2.6 1308 82.0
SMT-T [37] 12 2.4 638 82.2
RMT-T [15] 14 2.5 1438 82.4
SECViT-T 15 2.5 2004 82.7
Swin-T [41] 29 4.5 1723 81.3
PS-ViT-B14 [61] 21 5.4 1986 81.7
DVT-T2T-ViT-19 [52] 39 6.2 1268 81.9
SGFormer-S [17] 23 4.8 952 83.2
CMT-S [18] 25 4.0 846 83.5
CSwin-S [12] 35 6.9 972 83.6
SMT-S [37] 20 4.8 356 83.7
BiFormer-S [66] 26 4.5 766 83.8
SEC-Swin-T 29 4.8 1482 83.8
SECViT-S 27 4.6 998 84.3
Swin-S [41] 50 8.8 1006 83.0
SGFormer-M [17] 39 7.5 598 84.1
SMT-B [37] 32 7.7 237 84.3
BiFormer-B [66] 57 9.8 498 84.3
MaxViT-S [47] 69 11.7 546 84.5
CMT-B [18] 46 9.3 447 84.5
iFormer-B [44] 48 9.4 688 84.6
RMT-B [15] 54 9.7 430 85.0
SEC-Swin-S 50 9.2 804 85.0
SECViT-B 57 9.8 504 85.2
Swin-B [41] 88 15.5 768 83.5
CSWin-B [12] 78 15.0 660 84.2
SMT-L [37] 80 17.7 158 84.6
SGFormer-B [17] 78 15.6 388 84.7
iFormer-L [44] 87 14.0 410 84.8
MaxViT-B [47] 120 23.4 306 84.9
SEC-Swin-B 88 16.2 696 85.3
SECViT-L 101 18.2 398 85.7
Table 13: Comparison of models’ efficiency. Throughputs are measured on a single A100 with the batch size of 64.
Number of Vision Tokens Outputs by SEC.
MLLM is quite sensitive to the number of vision tokens. We conduct a detailed exploration based on LLaVA-1.5 regarding the number of vision tokens output by SEC, as shown in Tab. 14. The first row represents the speed and performance of the original LLaVA-1.5 without using SEC. Compared to LLaVA-1.5, employing SEC effectively reduces the number of vision tokens and improves inference efficiency. As the number of vision tokens decreases, the model’s performance shows a slight decline, but its efficiency is further enhanced.
V-T num Time Speed TextVQA GQA VQAv2 POPE MM-Vet
576+1 194s 1.0×1.0$\times$1.0 × 58.2 62.0 78.5 86.1 31.1
288+1 126s 1.5×1.5$\times$1.5 × 60.1(+1.9) 63.5(+1.5) 78.9(+0.4) 87.7(+1.6) 33.2(+2.1)
256+1 116s 1.7×1.7$\times$1.7 × 59.6(+1.4) 63.2(+0.3) 78.6(+0.1) 87.1(+1.0) 32.7(+1.6)
192+1 102s 1.9×1.9$\times$1.9 × 57.7(-0.5) 62.7(+0.7) 78.4(-0.1) 86.7(+0.6) 32.1(+1.0)
144+1 94s 2.1×2.1$\times$2.1 × 56.8(-1.4) 62.0(+0.0) 78.0(-0.5) 86.1(+0.0) 31.7(+0.6)
Table 14: Effect of the number of vision tokens outputs by SEC. “V-T num” means the number of vision tokens output by SEC. The experiments are conducted based on LLaVA-1.5 [39].
Different Methods for Merging Vision Tokens.
For MLLM, SEC uses an interleaved merge token approach to reduce the number of vision tokens. Conversely, we also explore a sequential merge token method to achieve a similar reduction. The comparison of these two methods is shown in Tab. 15. The direct sequential merge token approach may result in the loss of critical visual information, significantly degrading the model’s performance.
Method V-T num Time Speed TextVQA GQA VQAv2 POPE MM-Vet
Interleaved 288+1 14h 1.5×1.5$\times$1.5 × 60.1 63.5 78.9 87.7 33.2
Sequential 288+1 14h 1.5×1.5$\times$1.5 × 52.8(-7.3) 57.1(-6.2) 75.7(-3.2) 81.7(-6.0) 27.6(-5.6)
Interleaved 144+1 10h 2.1×2.1$\times$2.1 × 56.8 62.0 78.0 86.1 31.7
Sequential 144+1 10h 2.1×2.1$\times$2.1 × 47.2(-9.6) 53.6(-8.4) 71.7(-6.3) 80.0(-6.1) 22.3(-9.6)
Table 15: Different methods for merging vision tokens.
Appendix C More Clustering Results for Complex Scenes
To further illustrate the mechanism of SEC, we visualize more images in complex scenes and their clustering results, as shown in Fig. 6. Specifically, we visualize the clustering results of the first three stages of SECViT. The results further demonstrate that SEC can better learn fine-grained representations in the shallow layers of the model and semantic representations in the deeper layers.
Figure 6: Visualization results for complex scenes.